# Analyzing Similarity Metrics for Data Selection for Language Model Pretraining

### v2

@@ Abstract
Similarity between training examples is used to curate pretraining datasets for language models by many methods—for diversification and to select examples similar to high-quality data. However, similarity is typically measured with off-the-shelf embedding models that are generic or trained for tasks such as retrieval. This paper introduces a framework to analyze the suitability of embedding models specifically for data curation in the language model pretraining setting. We quantify the correlation between similarity in the embedding space to similarity in pretraining loss between different training examples, and how diversifying in the embedding space affects pretraining quality. We analyze a variety of embedding models in our framework, with experiments using the Pile dataset for pretraining a 1.7B parameter decoder-only language model. We find that the embedding models we consider are all useful for pretraining data curation. Moreover, a simple approach of averaging per-token embeddings proves to be surprisingly competitive with more sophisticated embedding models—likely because the latter are not designed specifically for pretraining data curation. Indeed, we believe our analysis and evaluation framework can serve as a foundation for the design of embedding models that specifically reason about similarity in pretraining datasets.
Machine Learning, ICML
@@ 1 Introduction
The recent success of general purpose language models (Brown et al., 2020; Chowdhery et al., 2023) is in no small part due to pretraining on large and diverse text corpora scraped from a variety of sources (Raffel et al., 2020; Gao et al., 2020; Penedo et al., 2024). Such pretraining allows the model to quickly adapt to various downstream tasks with just few-shot prompting or minimal finetuning. And so, researchers have explored a variety of approaches to assemble effective pretraining sets, typically by selecting a high-quality and diverse subset from larger corpora of examples scraped from various sources.
@@
These data curation approaches have shown promising results by improving example quality and by reducing redundancy in pretraining sets. Many of these methods use notions of similarity between examples. Similarity of an example to text from known high-quality sources (such as Wikipedia) has been used as a proxy for the quality of that example (Gunasekar et al., 2023; Penedo et al., 2024). Meanwhile, methods focused on diversification (Abbas et al., 2023; Tirumala et al., 2023) make direct use of similarity metrics to identify and remove redundant examples.
@@
In this context, similarity between examples has been measured in terms of distances in an embedding space, and these methods have typically used generic embeddings for this purpose—e.g., intermediate representations learned by language models, or those trained for semantic retrieval. But whether these embeddings are optimal for reasoning about similarity between pretraining examples—and whether there is any benefit to using with more sophisticated and potentially expensive models—remains an open question.
@@
Figure 1: Correlation between pretraining loss and embedding distance. Each row shows a pair of examples close in the embedding space (in particular, from the same cluster when using K-Means clustering in this space) with examples in different rows being relatively far from each other (from different clusters). We find that the close pairs of examples in each row tend to have similar pretraining losses, while there is a greater variation in losses across rows. Note that the close example pairs are ”thematically” similar, but have very different content. These results are from using the average embeddings from the final layer of a small decoder-only language model.
In this paper, we introduce a framework to analyze and quantify the suitability of a given embedding model for the task of data curation for pretraining language models. We consider the question of what a similarity metric should ideally capture for reasoning about pretraining examples. In this context, we characterize how well similarity in the embedding space correlates with similar pretraining behavior—in particular, similar pretraining losses with respect to the same model state (see Fig. 1). We also quantify whether the embedding space is able to separate examples from different data sources from among a mixture of these sources, with the assumption that these sources were curated by human expertise and combined since they were found to be complementary. And last but not least, we use the embeddings in a simple diversification recipe in curating a pretraining set and measure the gain in quality of the resulting pretrained model over a large number of downstream tasks.
@@
We conduct these experiments using the Pile (Gao et al., 2020) as our data corpus, and in the context of pretraining a 1.7B decoder-only language model with a UL2 objective (Tay et al., 2022). We evaluate a number of different embeddings models—from representations of generic language models (e.g., trained with mask-based reconstruction or mixtures of pretraining objectives) to embeddings specifically trained for retrieval or semantic equivalence.
Our results find that most embeddings bring some utility to the pretraining data curation setting—in particular, diversity-based pretraining data curation with all of the embeddings that we consider outperforms baseline training. But perhaps surprisingly, a very simple embedding approach—derived from averaging over the independent per-token embeddings from the input layer of a small language model—performs fairly competitively to more sophisticated and computationally expensive approaches.
These results also strongly suggest that embedding models for reasoning about pretraining examples should be specialized, and that the attributes that make an embedding suitable for semantic matching and retrieval may not translate to benefits for the pretraining setting. More broadly, our work establishes a framework to evaluate current embeddings, and facilitate the design of new ones, for the important task of data curation for language model pretraining.
@@ 2 Related Work
One family of works approaches data curation with the goal of retaining only high quality examples. The simplest approaches in this family look at heuristics that can filter out noisy or extremely low-quality instances such as documents containing less than 5 sentences or those with extremely short lines of text (Raffel et al., 2020; Xue, 2020). Beyond these heuristic approaches, a line of work focus on extracting high-quality examples based on similarity to known high-quality sources, such as Wikipedia or textbooks (Gunasekar et al., 2023; Penedo et al., 2024). Other works look at creating targeted sets—defining quality in terms of their relevance for particular downstream tasks (Xie et al., 2023), with some creating these by adaptively updating weights over domain mixtures (Xie et al., 2024; Jiang et al., 2024).
The other family of data curation approaches focus on pruning datasets to ensure a notion of diversity (Abbas et al., 2023, 2024; Tirumala et al., 2023). While some of these works (Abbas et al., 2024; Tirumala et al., 2023) compare different embedding models as part of brief ablations, these comparisons are in the narrow context of their specific divsersification algorithms, and do not provide intuition as to why one embedding may perform better than another.
Finally, it is worth mentioning that some works in data selection use the model being trained as part of the selection process, often through the use of influence functions or gradient-based methods (Garima et al., 2020; Xia et al., 2024; Engstrom et al., 2024). However, these are typically used during for data-curation for multi-epoch training on small datasets or only during the finetuning stage, and would be prohibitively expensive to use for pretraining.
@@
Another related line of work is the task of learning embedding models for text. Many various approaches to learn text embedding models, with objectives including mask-based reconstruction (Devlin, 2018; Liu et al., 2019), a combination of multiple different tasks (Cer, 2018), and contrastive learning-based approaches (Gao et al., 2021; Neelakantan et al., 2022; Izacard et al., 2022; Lee et al., 2024). More recent work has demonstrated extracting embeddings via prompting, specifically designed for predicting model performance (Sam et al., 2025).
These embedding models have largely been studied in the context for classification or similarity measures (Gomaa et al., 2013; Agirre et al., 2013, 2016), and recent focus has been improving performance on aggregate on large-scale benchmarks (Muennighoff et al., 2022) that are comprised of multiple tasks (e.g., retrieval, clustering, classification, ranking). Many models that achieve strong performance on these large-scale benchmarks have benefitted from scaling (Jiang et al., 2023; Chowdhery et al., 2023), with the help of distilled information from autoregressive models (Lee et al., 2024). However, this has conflicting incentives with their utility in the pretraining setting, as large text embedding models are impractical and too computationally expensive to run inference over the full pretraining corpus. To the best of our knowledge, our work provides the first study of various text embedding models for pretraining data curation.
3 Evaluating Embeddings for Pretraining Data Curation
We now describe our approach to analyze and evaluate a given embedding model in terms of its suitability for reasoning about similarity among pretraining examples. As mentioned above, similarity is used to (a) find examples that are similar to known “desirable” examples (e.g., examples of known high-quality or those representing downstream tasks of interest) and to (b) discover and remove redundancies in the pretraining corpus. Accordingly, we design experiments that measure an text embedding model’s performance towards these criteria.
@@ 3.1 Evaluating Correlation with Pretraining Loss
We begin by evaluating whether low distances in a model’s embedding space correlates with similar values of difficulty, which we measure by intermediate losses during language model pretraining. To do so, we first use a balanced K-Means clustering algorithm in the given embedding space to cluster all examples in the pretraining set for a target cluster size. Then, we look at the variance of loss values within each cluster. We note that we use balanced clustering (allowing only some variation in the sizes of different clusters) to ensure that average within-cluster variance is comparable across clusterings in different embedding spaces. We repeat this process for a multiple target cluster sizes for all embeddings.
When reporting our results, we measure variance reduction, or the ratio of the overall variance across all examples in the pretraining dataset to the variance in loss computed within clusters in the given embedding space. Specifically, we define variance reduction as
@@
where $C=\{C_{1},...,C_{m}\}$ is a partition of all datapoints in the dataset $D$, and $\mathbbm{C}$ is a uniform distribution over each disjoint set in the partition $C$. This metric can be interpreted as a measure of predicting pretraining generalization, as nearby points have similar loss values. We remark that random clusters achieve a variance reduction of 1, and larger values of variance reduction imply that points with more similar pretraining loss are clustered together.
A high value of variance reduction implies that similarity in the embedding space correlates well with pretraining loss. While an isolated pair of examples having similar pretraining loss may be entirely unrelated in quality and content, the fact that an embedding space that consistently brings together examples of similar loss values strongly suggests that these similar examples will behave similarly in terms of contributing to the quality of the language model.
@@
3.2 Measuring Cluster Purity of Data Sources
Next, we study the purity of the respective (K-Means) clusters generated with each considered embedding model with respect to the underlying data source or mixture. Large pretraining datasets such as the Pile (Gao et al., 2020) are comprised of various distinct yet complementary hand-curated data sources (e.g., high quality sources such as Wikipedia, code-based data, data from medical domains). While there may be some similarities between datapoints from different domains, we believe that clusters should generally group points from the same domain together and separate isntances from different domains. In fact, we remark that since embedding models are generally not trained with explicit knowledge of source labels or any domain metadata, and the ability to group together data from the same source and distinguish those from different sources shows an embedding model’s alignment with human judgement (or at least that of the dataset curators).
@@ 3.2 Measuring Cluster Purity of Data Sources
3.3 Diversification-based Pretraining Data Curation
Prior work has demonstrated that diversification, based on similarity in embedding spaces, can yield more effective pretraining datasets that in-turn lead to improved models (Tirumala et al., 2023). While the diversification techniques themselves tend to be sophisticated, our goal here is to evaluate the utility of a given embedding for diversification. Therefore, we use a simplified version of these diversification of approaches to (1) select a subset from a larger pretraining corpus, (2) pretrain a model on this curated subset, and then (3) report performance on a large set of downstream tasks.
Figure 2: Reduction in variance in loss among clusters (e.g., overall variance / clustering variance) as increase the number of gradient steps in pretraining (left). Reduction in variance as we vary the approximate cluster size used in K-means for clustering examples with various embedding models (right). We compute our results on the checkpoint after 26k gradient steps. In both figures, we compute our results over 50 million sampled clusters from the Pile, and we remark that larger values are better. The overall pretraining loss variances for checkpoints of 26k, 50k, and 76k training steps are 0.521, 0.508, and 0.489 respectively. We find that embeddings extracted from a small version of the downstream language model being trained best cluster points with similar pretraining losses, across different intermediate training checkpoints and cluster sizes.
We begin by clustering the larger pretraining corpus in the given embedding space; in contrast to Section 3.1 and Section 3.2, we do not use balanced clustering which can lead to clusters with a wide variation in “diameters” (distances among points in the same cluster). Instead, since the goal here is diversification, we find clusters such that distances between all pairs of points in the same cluster are within a specified threshold $\epsilon$.
Given that we select a single point from each cluster, this implies that we need the number of clusters to be equal or greater to the number of desired pretraining examples. In practice, we sweep a large grid of $\epsilon$ values and select the largest threshold that still produces a sufficient number of clusters. Another important point of note is that producing such a large number of clusters is computationally expensive. To be able to scale to such a large number of clusters, we use the reciprocal agglomerative clustering (RAC) (Sumengen et al., 2021) algorithm. This has the advantage of building a clustering from progressively larger values of $\epsilon$, which works well with our need of a grid search over values of $\epsilon$.
@@ 4 Experiments
We analyze various different approaches to extract text embeddings across multiple different pretraining settings.
@@ 4.1 Experiment Details
Text Embedding Models
@@
For text embedding models, we compare Universal Sentence Encoders (USE) (Cer, 2018), which are general purpose text encoders, and Gecko (Lee et al., 2024), a retrieval-focused text embedding model trained via synthetic data distilled from LLMs, and BERT (Devlin, 2018) models. We also compare against various techniques to extract embeddings from a small version of the downstream language model we are training (e.g., a  200M parameter language model). First, we extract an embedding by simply averaging the token embeddings matrix over all tokens in the input sequence, which we refer to as LM Token Embeds. We note that this is equivalent to a learned unigram model (i.e., ignoring all positional information). We also extract an embeddings from a forward pass of the language model by averaging the activations over all tokens in the input sequence at the final hidden layer, which we refer to as LM Output Embeds. In our diversity-based pretraining data curation experiments, we also add a comparison to a naive, random subset selection (Baseline) to evaluate how much benefit is added from our standardized curation strategy with these various embedding models.
@@
To make clustering with a large number of output clusters feasible and efficient at pretraining data scales, we perform dimensionality reduction using PCA on a large subset of the datapoints ($\sim$500,000 examples) and extract the top 64 dimensions upon which to project. The projected embeddings are then normalized to have a unit L2 norm. We also explore using random projections (Bingham & Mannila, 2001) as a different dimensionality reduction strategy, finding that random projections do not perform as well as PCA (see Section 4.5 for ablation studies).
Dataset Details
For all of our experiments (e.g., pretraining data curation and predicting loss generalization), we use the Pile (Gao et al., 2020). We pack together documents into a sequence of length 1280, with “[eod]” as delimeters between documents.
Pretraining Details
@@
For our pretraining experiments, we train a 1.7B parameter decoder-only language model with a UL2 objective (Tay et al., 2022), on the curated subset of the Pile data through the process outlined in Section 3.3. We use a selection budget of 200B tokens, which is approximately 20% of sequences from the Pile. This corresponds to roughly 170 million clusters for each embedding model, with an average cluster size of 5 examples.
@@
We pretrain with a learning rate of 0.001 with a linear decay and a batch size of 1024. For our tokenizer, we use sentencepiece with a vocabulary size of 256k tokens. For our pretraining evaluation, we consider the same downstream evaluation datasets from the work of Brown et al. (2020). This invovles a wide variety of 1-shot scoring tasks, as well as open-ended text generation tasks. The full details about our evaluation sets are deferred to Section B.1. For the embeddings produced via LM Token Embeds and LM Output Embeds, we train a 200M parameter language model in the same fashion as above, where we train on a total of 200B randomly selected tokens.

### v3

@@ Abstract
Measuring similarity between training examples is critical for curating high-quality and diverse pretraining datasets for language models. However, similarity is typically computed with a generic off-the-shelf embedding model that has been trained for tasks such as retrieval. Whether these embedding-based similarity metrics are well-suited for pretraining data selection remains largely unexplored. In this paper, we propose a new framework to assess the suitability of a similarity metric specifically for data curation in language model pretraining applications. Our framework’s first evaluation criterion captures how well distances reflect generalization in pretraining loss between different training examples. Next, we use each embedding model to guide a standard diversity-based data curation algorithm and measure its utility by pretraining a language model on the selected data and evaluating downstream task performance. Finally, we evaluate the capabilities of embeddings to distinguish between examples from different data sources. With these evaluations, we demonstrate that standard off-the-shelf embedding models are not well-suited for the pretraining data curation setting, underperforming even remarkably simple embeddings that are extracted from models trained on the same pretraining corpus. Our experiments are performed on the Pile, for pretraining a 1.7B parameter language model on 200B tokens. We believe our analysis and evaluation framework serves as a foundation for the future design of embeddings that specifically reason about similarity in pretraining datasets.
@@ 1 Introduction
The recent success of language models (Brown et al., 2020; Chowdhery et al., 2023) is in no small part due to pretraining on large and diverse text corpora scraped from a variety of sources (Raffel et al., 2020; Gao et al., 2020; Penedo et al., 2024). Researchers have explored a variety of approaches to assemble effective pretraining sets, typically by selecting a high-quality and diverse subset from a larger corpus of examples scraped from multiple data sources. These data curation approaches have shown promising results by improving example quality and reducing redundancy in pretraining sets. Many of these methods use notions of similarity between examples. Similarity of an example to text from known high-quality sources (such as Wikipedia) has been used as a proxy for the quality of that example (Gunasekar et al., 2023; Penedo et al., 2024). Meanwhile, methods focused on diversification (Abbas et al., 2023; Tirumala et al., 2023) make direct use of similarity metrics to identify and remove redundant examples.
@@
In this context, similarity between training examples has often been measured in terms of distances in an embedding space. Many approaches have typically used generic embeddings for this purpose (Chang and Jia, 2023; Vo et al., )—off-the-shelf embedding models that have been trained for tasks such as semantic retrieval or mask-based reconstruction. However, whether these embeddings are optimal for reasoning about similarity between pretraining examples and whether there is any benefit to using more sophisticated and computationally expensive models remains an open question. Furthermore, evaluating such embeddings in pretraining applications is fundamentally challenging, as the sheer scale of pretraining data makes many experiments intractable. Such challenges necessitate specialized evaluations — new criteria for assessing the utility of embeddings in pretraining applications.
In this paper, we propose a novel evaluation framework to assess the suitability of an embedding model for curating pretraining data for language models. Our goal is to establish a new standard for evaluating similarity metrics in the context of language model pretraining. We begin by asking: what should a similarity metric ideally capture to be useful in this setting? Unlike prior work that evaluates embeddings for downstream tasks like retrieval or classification, we argue that embeddings used for data curation must reflect training dynamics—such as generalization behavior or corpus redundancy—that are unique to the pretraining setting.
@@
Figure 1: A visualization of the correlation between pretraining loss and embedding distance. Each row shows a pair of examples close in embedding space (from the same K-means cluster), with examples in different rows being far from each other (from different clusters). We find that close pairs of examples tend to have similar pretraining losses, while there is a greater variation in losses across clusters. Close example pairs are "thematically" similar but have different content. These results are from averaged embeddings from the final layer of a small decoder-only language model.
Our first evaluation criterion measures how well distances in the embedding space correlate with generalization behavior during pretraining—specifically, whether nearby examples in embedding space exhibit similar pretraining loss under the same model state (see Fig. 1). This captures their utility of similarity metrics in loss-based data selection strategies (Jiang et al., 2023) or for selecting examples similar to high-quality references (Penedo et al., 2024). Next, we evaluate the embedding model’s utility in guiding a simple diversity-based curation method (similar to those employed in (Tirumala et al., 2023)) and assess its effectiveness by pretraining a model on the curated subset and measuring downstream task performance. Finally, we also measure whether the embedding space can distinguish between examples from different data sources. While not directly used in data selection, this serves as a useful proxy for embedding quality under the assumption that these sources were curated by human expertise and reflect meaningful human-curated structure.
@@
We conduct all of our experiments using the Pile (Gao et al., 2020) as our data corpus, and in the context of pretraining a 1.7B parameter decoder-only language model with a UL2 objective (Tay et al., 2022) on 200B tokens. We evaluate a number of different off-the-shelf embedding models—from representations of generic language models (e.g., trained with mask-based reconstruction or mixtures of pretraining objectives) to embeddings specifically trained for retrieval or semantic equivalence. We compare these to two specialized approaches that are derived from smaller, proxy models of the downstream model being trained on the same pretraining set: (1) computing the average token embedding at the final hidden state, and (2) a significantly more computationally efficient and simple average of the token-embeddings from the input layer (i.e., requiring no forward passes).
Despite their strong performance on standard semantic similarity and retrieval benchmarks, we find that off-the-shelf embedding models underperform in the context of pretraining data curation, suggesting that existing benchmarks may not reflect the inductive biases most relevant to language model pretraining. Surprisingly, even this simple method that averages input token embeddings—requiring no forward pass—matches or outperforms these more computationally intensive models. Adding a forward pass in this specialized model yields further improvements. These results strongly suggest that embedding models for reasoning about similarity between pretraining examples should be specialized to the data distribution at hand, and that properties that make embeddings suitable for retrieval and semantic matching may not transfer to the pretraining setting. More broadly, our work establishes a standardized framework to evaluate and highlight failures of current off-the-shelf models, facilitating the design of new models tailored for data curation in language model pretraining.
@@ 2 Related Work
One family of works approaches data curation with the goal of retaining only high quality examples. The simplest approaches in this family look at heuristics that can filter out noisy or extremely low-quality instances, such as documents containing less than 5 sentences or those with extremely short lines of text (Raffel et al., 2020; Xue, 2020). Beyond these heuristic approaches, a line of work focuses on extracting high-quality examples based on similarity to known high-quality sources, such as Wikipedia or textbooks (Gunasekar et al., 2023; Penedo et al., 2024). Other works look at creating targeted sets—defining quality in terms of their relevance for particular downstream tasks (Xie et al., 2023), with some creating these by adaptively updating weights over domain mixtures (Xie et al., 2024; Jiang et al., 2024). The other family of data curation approaches focus on pruning datasets to ensure a notion of diversity and reducing redundancies in pretraining examples (Abbas et al., 2023, 2024; Tirumala et al., 2023). While some of these works (Abbas et al., 2024; Tirumala et al., 2023) compare different embedding models as part of brief ablations, these comparisons are in the narrow context of their specific diversification algorithms. Finally, it is worth mentioning that some works in data selection use the model being trained as part of the selection process, often through the use of influence functions (Garima et al., 2020; Xia et al., 2024; Engstrom et al., 2024). However, these are typically only used in data-curation on small datasets or only during finetuning, and would be prohibitively expensive in pretraining.
@@
Another related line of work is the task of learning embedding models for text. Many various approaches to learn text embedding models, with objectives including mask-based reconstruction (Devlin, 2018; Liu et al., 2019), a combination of multiple different tasks (Cer, 2018), and contrastive learning-based approaches (Gao et al., 2021; Neelakantan et al., 2022; Izacard et al., 2022; Lee et al., 2024). More recent work has studied extracting embeddings from standard decoder-only language models (BehnamGhader et al., 2024), even via prompting (Sam et al., 2025).
These embedding models have largely been studied in the context for classification or similarity measures (Gomaa et al., 2013; Agirre et al., 2013, 2016), and recent focus has been on improving performance on aggregate on large-scale benchmarks (Muennighoff et al., 2022) that are comprised of multiple tasks (e.g., retrieval, clustering, classification, ranking). Many models that achieve strong performance on these large-scale benchmarks have benefitted from scaling (Jiang et al., 2023; Chowdhery et al., 2023), with the help of distilled information from autoregressive models (Lee et al., 2024). However, this has conflicting incentives with their utility in the pretraining setting, as large text embedding models are impractical and too computationally expensive to run inference over the full pretraining corpus. To the best of our knowledge, our work provides the first study of various text embedding models for pretraining data curation.
3 An Evaluation Framework for Embeddings in Pretraining Data Curation
We now describe our new framework for analyzing text embedding models in terms of their suitability for reasoning about similarity among pretraining examples. As previously mentioned, similarity is used to (a) find examples that are similar to known “desirable” examples (e.g., examples of known high-quality or those representing downstream tasks of interest) and to (b) discover and remove redundancies in the pretraining corpus. Accordingly, we design standardized experiments that measure a text embedding model’s performance towards these criteria.
@@ 3.1 Evaluating Correlation with Pretraining Loss
We begin by evaluating whether low distances in a model’s embedding space correlate with similar values of difficulty, which we measure by intermediate losses during language model pretraining. Examples of positive correlations are visualized in Figure˜1. To do so, we first use a balanced K-Means clustering algorithm in the embedding space to cluster all examples in the pretraining set for a target cluster size. Then, we look at the variance of loss values within each cluster. This within-cluster variance captures whether points with similar pretraining loss (or difficulty) are grouped together and are close in embedding space. Crucially, we note that the converse is not necessarily true; two very different examples can have the same loss because they happen to be equally difficult to learn. This asymmetry is an important reason we chose our cluster-based approach rather than alternatives such as pairwise correlation. We note that we use balanced clustering (allowing only some variation in the sizes of different clusters) to ensure that average within-cluster variance is comparable across clusterings from different embedding models. We repeat this process for multiple target cluster sizes for all embeddings.
When reporting our results, we measure variance reduction, or the ratio of the overall variance across all examples in the pretraining dataset to the variance in loss computed within clusters in the given embedding space. Formally, we define variance reduction as
@@
where $C=\{C_{1},...,C_{m}\}$ is a clustering or partition of all datapoints in the dataset $D$, and $\mathbbm{C}$ is a uniform distribution over each disjoint set in the partition $C$. Random clustering achieves a variance reduction of 1, and larger values of variance reduction imply that points with more similar pretraining loss are clustered together.
A high value of variance reduction implies that similarity in the embedding space correlates well with pretraining loss. While an isolated pair of examples having similar pretraining loss may be entirely unrelated in quality and content, the fact that an embedding space that consistently brings together examples of similar loss values strongly suggests that these similar examples will behave similarly in terms of contributing to the quality of the language model (when used in pretraining).
@@
3.2 Diversification-based Pretraining Data Curation
Prior work has demonstrated that data curation schemes that encourage diversification, based on similarity in embedding spaces, leads to improved models when trained on the curated subsets (Abbas et al., 2023; Tirumala et al., 2023). While the diversification techniques tend to be sophisticated, our goal here is to evaluate the utility of a given embedding for diversification. Therefore, we use a simplified version of these approaches to (1) select a subset from a pretraining corpus, (2) pretrain a model on this curated subset, and then (3) report performance on a large set of downstream tasks.
We begin by clustering the larger pretraining corpus in the given embedding space; in contrast to Section˜3.1, we do not use balanced clustering, which can lead to clusters with a wide variation in “diameters” (distances among points in the same cluster). Instead, since the goal here is diversification, we find clusters such that distances between all pairs of points in the same cluster are within a specified threshold $\epsilon$.
Given that we select a single point from each cluster, this implies that we need the number of clusters to be equal to or greater than the number of desired pretraining examples. In practice, we sweep a large grid of $\epsilon$ values and select the largest threshold that still produces a sufficient number of clusters. Another important point of note is that producing such a large number of clusters is computationally expensive. To be able to scale to such a large number of clusters, we use the reciprocal agglomerative clustering (RAC) (Sumengen et al., 2021) algorithm. This has the advantage of building a clustering from progressively larger values of $\epsilon$, which works well with our need for a grid of values of $\epsilon$.
3.3 Measuring Cluster Purity with Respect to Data Sources
Large pretraining datasets such as the Pile (Gao et al., 2020) are comprised of various distinct yet complementary hand-curated data sources (e.g., high quality sources such as Wikipedia, code-based data, data from medical domains). While there may be some similarities between datapoints from different domains, we believe that clusters should generally group points from the same domain together and separate instances from different domains. In fact, since embedding models are generally not trained with explicit knowledge of source labels or any domain metadata, the ability to group together data from the same source and distinguish those from different sources shows an embedding model’s alignment with meaningful human-curated structure (or at least that of the dataset curators). This also serves as a useful proxy of embedding model capacity, as they should contain sufficient information to well-separate these different sources.
@@ 4 Experiments
We demonstrate the utility of our framework in evaluating off-the-shelf embedding models in measuring pretraining similarity and compare them with simple ways to produce specialized embeddings.
@@ 4.1 Experiment Details
Embedding Models
For the off-the-shelf embedding models, we consider: (1) Universal Sentence Encoders (USE) (Cer, 2018), which are general-purpose text encoders trained on a combination of objectives, (2) Gecko (Lee et al., 2024), a retrieval-focused text embedding model trained via synthetic data distilled from LLMs, and (3) BERT (Devlin, 2018) embedding models. These encapsulate a variety of different training objectives and reflect common embedding model choices in the field.
@@
We also consider a few standard approaches to extract specialized embeddings that are extracted from a small version of the downstream language model we are training (e.g., a  200M parameter language model). The first and most simple approach is to extract an embedding by simply averaging the token embeddings matrix over all tokens in the input sequence, which we refer to as LM Token Embeds. This is extremely efficient as it only requires looking up the token embedding matrix and does not involve any forward passes. We note that this is also equivalent to a learned unigram model (i.e., ignoring all positional information). The second approach is to extract an embedding from the forward pass of the language model by averaging the activations over all tokens in the input sequence at the final hidden layer, which we refer to as LM Output Embeds.
Finally, in our diversity-based pretraining data curation experiments, we also add a comparison to a naive, random subset selection (Baseline). This evaluates how much benefit is observed from our standardized curation strategy with these various embedding models over standard baseline training.
Clustering Details
@@
To make clustering with a large number of output clusters feasible and efficient at pretraining data scales, we perform dimensionality reduction using PCA on a large subset of the datapoints ($\sim$500,000 examples) and extract the top 64 dimensions upon which to project. The projected embeddings are then normalized to have a unit L2 norm. We ablate on this choice for dimensionality reductions by comparing with random projections (Bingham and Mannila, 2001) as an alternative dimensionality reduction strategy; we find that random projections do not perform as well as PCA (Section˜4.5). Other work (Tirumala et al., 2023) tackles this computational challenge by running clustering only with a much smaller subset of data, rather than performing dimensionality reduction and then clustering over all the data.
Figure 2: Variance reduction as we vary average cluster size. Larger values are better. Results are computed over 50 million sampled clusters from the Pile, where pretraining losses are computed after 26k gradient steps. Specialized embeddings yield higher variance reduction than off-the-shelf models for all cluster sizes.
Figure 3: Variance reduction as we increase the number of gradient steps in pretraining. Larger values are better. Results are computed over 50 million sampled clusters from the Pile with an average cluster size of 50. Benefits in variance reduction remain consistent throughout pretraining.
Dataset and Pretraining Details
@@
For all of our experiments (e.g., pretraining data curation and predicting loss generalization), we use the Pile (Gao et al., 2020). We pack together documents into a sequence of length 1280, with “[eod]” as delimiters between documents. For our pretraining experiments, we train a 1.7B parameter decoder-only language model with a UL2 objective (Tay et al., 2022). We curated pretraining subsets from the Pile through the process outlined in Section˜3.2. We use a selection budget of 200B tokens, or approximately 20% of the Pile. This corresponds to roughly 170 million clusters for each embedding model, with an average cluster size of 5 examples.
@@
We pretrain with a learning rate of 0.001 with a linear decay and a batch size of 1024. For our tokenizer, we use sentencepiece with a vocabulary size of 256k tokens. For our pretraining evaluation, we consider the set of 23 downstream evaluation datasets from the work of Brown et al. (2020). This involves a wide variety of 1-shot scoring tasks, as well as open-ended text generation tasks. More details about our evaluation sets are deferred to Section˜B.1. For the embeddings produced via LM Token Embeds and LM Output Embeds, we train a 200M parameter language model in the same fashion as above, where we also train on a total of 200B randomly selected tokens.
Table 1: Average downstream task performance of embedding models in the diversity-based curation schemes of a 200B token subset from the Pile, to pretrain a 1.7B parameter decoder-only language model. Bolding and italicizing denote the best and second-best performing methods on each task, respectively. Results are averaged over 3 pretraining runs, and average results are mean $\pm$ standard error. Specialized embeddings (i.e., extracted from a small version of the downstream language model) and Gecko perform the best for diversity-based pretraining data curation.
Task Baseline USE Gecko BERT LM Token Embeds LM Output Embeds
ARC - Challenge 32.4 33.0 33.7 32.7 32.5 33.5
ARC - Easy 63.8 65.1 64.2 65.1 64.9 65.5
BoolQ 56.5 59.2 58.3 60.1 62.9 61.9
SuperGLUE - CB 42.4 43.5 48.2 41.7 42.9 48.8
SuperGLUE - Copa 75.3 76.3 77.3 74.0 74.7 78.0
HellaSwag 55.0 56.7 57.1 56.5 56.8 57.5
Multi RC 57.7 55.7 55.9 56.3 56.0 55.2
OpenBook QA 46.3 46.2 46.3 46.6 45.9 46.7
PiQA 72.3 73.0 72.7 73.6 72.8 73.6
Race H 38.0 37.9 38.5 38.8 38.8 38.7
Race M 51.7 52.0 53.3 52.9 52.1 52.6
ReCoRD 85.0 84.4 84.6 84.7 84.5 84.9
RTE 54.5 54.9 51.5 54.9 51.6 52.8
Story Cloze 73.9 74.1 74.3 74.0 73.3 74.5
WiC 48.2 47.9 47.5 47.5 48.2 47.3
Winograd 74.0 76.0 77.1 77.3 77.1 76.6
WinoGrande 59.2 58.9 59.0 59.5 60.0 59.3
WSC 74.1 73.9 73.7 74.7 73.2 75.0
Lambada 21.4 29.9 40.1 34.7 31.4 33.8
Natural Questions 10.0 10.8 10.1 10.5 11.1 11.4
SQuAD v2 51.1 54.3 51.7 51.8 58.6 55.4
TriviaQA Wiki 34.4 35.0 33.0 35.0 34.3 36.1
Web Questions 17.1 17.0 17.9 18.0 18.7 19.1
Average 51.9 $\pm$ 0.1 52.9 $\pm$0.1 53.3 $\pm$ 0.3 53.1 $\pm$ 0.2 53.1 $\pm$ 0.3 53.8 $\pm$ 0.1